OpenStudy (anonymous):

Select the correct reasons for statements 1 and 3 to complete the proof. Prove: 5x + 9 • x • 3 = 32x 1. 5x + 9 • x • 3 = 5x + 9 • 3 • x 1. ? 2. 5x + 9 • x • 3 = 5x + 27x 2. Multiplication 3. 5x + 9 • x • 3 = 32x 3. ? A. Associative Property of Addition, Addition B. Associative Property of Multiplication, Addition C. Commutative Property of Multiplication, Addition D. Commutative Property of Multiplication, Multiplication

OpenStudy (agent47):

What has changed in step 1? 9*x*3 became 9*3*x. What property is that?

OpenStudy (anonymous):

dont know

OpenStudy (agent47):

a*b=b*a, what property is that?

OpenStudy (anonymous):

commutative

OpenStudy (agent47):

yes, and what are we doing with the terms?

OpenStudy (anonymous):

i don't know

OpenStudy (agent47):

a * b <-- I am ________ a by b. What's the missing word, come on.

OpenStudy (anonymous):

..... I still don't know

OpenStudy (agent47):

4*2=8 What did I just do?

OpenStudy (anonymous):

multiplacation

OpenStudy (agent47):

YES

OpenStudy (agent47):

So first step is commutative property of multiplication.

OpenStudy (anonymous):

ok

OpenStudy (agent47):

5x + 27x = 32x what did I do here?

OpenStudy (anonymous):

addition

OpenStudy (agent47):

so what is the answer?

OpenStudy (anonymous):

C.

OpenStudy (anonymous):

i understand it now

OpenStudy (agent47):

good

OpenStudy (anonymous):

thank you

OpenStudy (agent47):

np